. A couple plans to have two children. Assume that on any given birth, having a boy and having a girl are equally likely.
(a) What is the probability that exactly one of the children is a boy?
(b) What is the probability that at least one of the children is a boy?


Answer by stanbon(75887)

A couple plans to have two children. Assume that on any given birth, having a boy and having a girl are equally likely.
---
Binomial with n=2 ; p = 1/2 ; q = 1/2
----------------------------------------------------
(a) What is the probability that exactly one of the children is a boy?
P(one boy) = P(boy first then girl) + P(girl first then boy)
= (1/2)(1/2) + (1/2)(1/2)
= 1/4 + 1/4
= 1/2
---------------------
(b) What is the probability that at least one of the children is a boy?
P(one boy) + P(2 boys) = 1/2 + (1/2)(1/2) = 3/4
====================================================
Cheers,
stan H.